# Joint parameter estimations for spin glasses

### v1

@@ 2.1. Auxiliary Lemmas
To prove the Theorem 1.4, we first establish the following universality result in Lemma 3.1, which shows that the free energies associated with the subgaussian matrix $G=(g_{ij})_{1\leqslant i,j\leqslant N}$ is close to that associated to the standard Gaussian disorder. For any $x=(x_{e})_{e\in E}$, set
$F_{x}(\beta,h):=\frac{1}{N}\ln\sum_{{\sigma}\in\{-1,+1\}^{N}}e^{\beta H_{x}({\sigma})+h\sum_{i=1}^{N}\sigma_{i}},$
where
$H_{x}({\sigma}):=\frac{1}{\sqrt{d}}\sum_{(i,j)\in E}x_{ij}\sigma_{i}\sigma_{j}.$
Note that $F_{g}(\beta,h)=F_{N}(\beta,h).$ Let $z=(z_{e})_{e\in E}$ be i.i.d. standard normal independent of $g.$
Lemma 3.1.
There exists a universal constant $C$ independent of $N,d,\beta,h$ such that
$\left|\operatorname{\mathds{E}}F_{g}(\beta,h)-\operatorname{\mathds{E}}F_{z}(\beta,h)\right|$ $\leqslant\frac{C\beta^{3}}{\sqrt{d}}.$
Proof.
For $0\leqslant t\leqslant 1$, define
$\phi(t)=\frac{1}{N}\operatorname{\mathds{E}}\ln e^{\beta H_{N,t}({\sigma})},$
where
$H_{N,t}({\sigma})$ $:=\sqrt{t}H_{g}({\sigma})+\sqrt{1-t}H_{z}({\sigma})+h\sum_{i=1}^{N}\sigma_{i}.$
Then
$\phi^{\prime}(t)$ $=\frac{\beta}{2N}\Bigl{(}\frac{\operatorname{\mathds{E}}\langle H_{g}({\sigma})\rangle_{t}}{\sqrt{t}}-\frac{\operatorname{\mathds{E}}\langle H_{z}({\sigma})\rangle_{t}}{\sqrt{1-t}}\Bigr{)}=\frac{\beta}{2N\sqrt{d}}\operatorname{\mathds{E}}\Bigl{[}\sum_{(i,j)\in E}\frac{\langle g_{ij}\sigma_{i}\sigma_{j}\rangle_{t}}{\sqrt{t}}-\sum_{(i,j)\in E}\frac{\langle z_{ij}\sigma_{i}\sigma_{j}\rangle_{t}}{\sqrt{1-t}}\Bigr{]}.$
Here from the Gaussian integration by parts in the Gaussian disorder,
$\frac{\operatorname{\mathds{E}}\langle z_{ij}\sigma_{i}\sigma_{j}\rangle_{t}}{\sqrt{1-t}}$ $=\frac{\beta}{\sqrt{d}}\operatorname{\mathds{E}}\bigl{\langle}1-\sigma_{i}^{1}\sigma_{j}^{1}\sigma_{i}^{2}\sigma_{j}^{2}\bigr{\rangle}_{t}.$
On the other hand, from the approximate Gaussian integration by parts in the disorder $g$ (see, for example, [carmona2006universality]), there exists an absolute constant $C$ independent of anything else such that
$\Bigl{|}\frac{\operatorname{\mathds{E}}\langle g_{ij}\sigma_{i}\sigma_{j}\rangle_{t}}{\sqrt{t}}-\frac{\beta}{\sqrt{d}}\operatorname{\mathds{E}}\bigl{\langle}1-\sigma_{i}^{1}\sigma_{j}^{1}\sigma_{i}^{2}\sigma_{j}^{2}\bigr{\rangle}_{t}\Bigr{|}\leqslant\frac{C\beta^{2}}{d}.$
Consequently,
$|\phi^{\prime}(t)|$ $\leqslant\frac{C\beta^{3}|E|}{2d^{3/2}N}=\frac{C\beta^{3}}{\sqrt{d}}$
and this completes our proof. $\blacksquare$
We also establish the following concentration result for the free energy $F_{N}(\beta,h)$ for later use.
Proposition 3.2.
Assume that $(g_{e})_{e\in E}$ is a collection of independent $\kappa^{2}$-subgaussian random variables. For $t>0$, we have
$\mathbb{P}\bigl{(}|F_{N}(\beta,h)-\operatorname{\mathds{E}}F_{N}(\beta,h)|\geqslant t\bigr{)}$ $\leqslant 2\exp\Bigl{(}-\min\Bigl{(}\frac{(\ln 2)Nt^{2}}{2\beta^{2}\kappa^{2}},\frac{\sqrt{d\ln 2}Nt}{2\beta\kappa}\Bigr{)}\Bigr{)}.$
Proof.
Denote $f_{N}(\beta,h):=NF_{N}(\beta,h).$ Let $(g_{i})_{1\leqslant i\leqslant|E|}$ be an enumeration of $(g_{e})_{e\in E}.$ We introduce the following filtration $\mathcal{F}_{k}=\sigma(g_{k}:1\leqslant k\leqslant|E|)$ for $1\leqslant k\leqslant|E|$ and $\mathcal{F}_{0}=\emptyset.$ For each $1\leqslant k\leqslant|E|,$ define the martingale difference
$X_{k}$ $:=\operatorname{\mathds{E}}\bigl{[}f_{N}(\beta,h)\big{|}\mathcal{F}_{k}\bigr{]}-\operatorname{\mathds{E}}\bigl{[}f_{N}(\beta,h)\big{|}\mathcal{F}_{k-1}\bigr{]}.$
Then $\operatorname{\mathds{E}}[X_{k}|\mathcal{F}_{k-1}]=0$ and
$\sum_{k=1}^{|E|}X_{k}$ $=f_{N}(\beta,h)-\operatorname{\mathds{E}}\bigl{[}f_{N}(\beta,h)\bigr{]}.$
Fix $k$. Let $\tilde{f}_{N}(\beta,h)$ be the same as $f_{N}(\beta,h)$ with $g_{k}=0$ and $\langle\cdot\rangle_{k}$ is the corresponding Gibbs expectation. Then
$X_{k}$ $=\operatorname{\mathds{E}}\bigl{[}f_{N}(\beta,h)-\tilde{f}_{N}(\beta,h)\big{|}\mathcal{F}_{k}\bigr{]}-\operatorname{\mathds{E}}\bigl{[}f_{N}(\beta,h)-\tilde{f}_{N}(\beta,h)\big{|}\mathcal{F}_{k-1}\bigr{]}$
$=\operatorname{\mathds{E}}\bigl{[}\ln\bigl{\langle}\exp\bigl{(}\beta g_{k}\sigma_{e}/\sqrt{d}\bigr{)}\bigr{\rangle}_{k}\big{|}\mathcal{F}_{k}\bigr{]}-\operatorname{\mathds{E}}\bigl{[}\ln\bigl{\langle}\exp\bigl{(}\beta g_{k}\sigma_{e}/\sqrt{d}\bigr{)}\bigr{\rangle}_{k}\big{|}\mathcal{F}_{k-1}\bigr{]}$
$\leqslant\beta d^{-1/2}\operatorname{\mathds{E}}\bigl{[}|g_{k}|\big{|}\mathcal{F}_{k}\bigr{]}+\beta d^{-1/2}\operatorname{\mathds{E}}\bigl{[}|g_{k}|\big{|}\mathcal{F}_{k-1}\bigr{]}$
$\leqslant\frac{\beta|g_{k}|}{\sqrt{d}}+\frac{\beta\operatorname{\mathds{E}}|g_{k}|}{\sqrt{d}},$
where $\sigma_{e}=\sigma_{i}\sigma_{j}$ and $e=(i,j)$ is the $k$-th edge in the enumeration. From the fact that $g_{k}$ is $\kappa^{2}$-subgaussian and Jensen’s inequality, we have
$\operatorname{\mathds{E}}\bigl{[}e^{\frac{|X_{k}|}{A}}\big{|}\mathcal{F}_{k-1}\bigr{]}$ $\leqslant\operatorname{\mathds{E}}\exp\left(\frac{\beta|g_{k}|}{A\sqrt{d}}+\frac{\beta\operatorname{\mathds{E}}|g_{k}|}{A\sqrt{d}}\right)$
$\leqslant\exp\left(\frac{\beta^{2}\kappa^{2}}{2A^{2}d}+\frac{\beta\operatorname{\mathds{E}}|g_{k}|}{A\sqrt{d}}\right)$
$\leqslant\exp\left(\frac{\beta^{2}\kappa^{2}}{2A^{2}d}\right)\operatorname{\mathds{E}}\exp\left(\frac{\beta|g_{k}|}{A\sqrt{d}}\right)=e^{\frac{\beta^{2}\kappa^{2}}{A^{2}d}}=e^{\ln 2}=2,$
provided
$A^{2}\equiv\frac{1}{\ln 2}\cdot\frac{\beta^{2}\kappa^{2}}{d}.$
Hence, from Bernstein’s inequality,
$\mathbb{P}\Bigl{(}\sum_{k=1}^{|E|}X_{k}\geqslant t\Bigr{)}$ $\leqslant\exp\Bigl{(}-\min\Bigl{(}\frac{t^{2}}{4|E|A^{2}},\frac{t}{2A}\Bigr{)}\Bigr{)}.$
and thus, using $|E|=2^{-1}\sum_{i\in[N]}d_{i}=2^{-1}Nd,$
$\mathbb{P}\left(\sum_{k=1}^{|E|}X_{k}\geqslant Nt\right)$ $\leqslant\exp\Bigl{(}-\min\Bigl{(}\frac{N^{2}t^{2}}{4|E|A^{2}},\frac{Nt}{2A}\Bigr{)}\Bigr{)}$
$=\exp\Bigl{(}-\min\Bigl{(}\frac{(\ln 2)dN^{2}t^{2}}{4\beta^{2}\kappa^{2}|E|},\frac{\sqrt{d\ln 2}Nt}{2\beta\kappa}\Bigr{)}\Bigr{)}$
$=\exp\Bigl{(}-\min\Bigl{(}\frac{(\ln 2)Nt^{2}}{2\beta^{2}\kappa^{2}},\frac{\sqrt{d\ln 2}Nt}{2\beta\kappa}\Bigr{)}\Bigr{)}.$
$\blacksquare$
Proof of Theorem 1.4.
From the Gaussian integration by parts, we have
$\partial_{\beta}\operatorname{\mathds{E}}F_{z}(\beta,h)$ $=\frac{\beta}{N}\operatorname{\mathds{E}}\bigl{\langle}H_{z}({\sigma})\bigr{\rangle}$
$=\frac{\beta^{2}}{Nd}\Bigl{(}|E|-\sum_{(i,j)\in E}\operatorname{\mathds{E}}\langle\sigma_{i}^{1}\sigma_{j}^{1}\sigma_{i}^{2}\sigma_{j}^{2}\rangle\Bigr{)}=\frac{\beta^{2}}{2}\Bigl{(}1-\frac{2}{Nd}\sum_{(i,j)\in E}\operatorname{\mathds{E}}\langle\sigma_{i}^{1}\sigma_{j}^{1}\sigma_{i}^{2}\sigma_{j}^{2}\rangle\Bigr{)},$
[…]
$U=\big{\{}i:|m_{i}({\sigma})|\leqslant M,\ d_{i}\geqslant d/4\big{\}},$
then w.h.p., $|U|\geqslant cn$.
Notice that
$\langle\sigma_{i}\sigma_{j}\rangle$ $=\bigl{\langle}\tanh(\beta m_{i}({\sigma})+h)\sigma_{j}\bigr{\rangle}.$
We have
$\langle\sigma_{i}\sigma_{j}\rangle$ $=\langle\tanh(\beta m_{i}({\sigma})+h)\sigma_{j};i\in U\rangle+\langle\tanh(\beta m_{i}({\sigma})+h)\sigma_{j};i\in U^{c}\rangle$
$\leqslant\tanh(\beta M+h)\bigl{\langle}I(i\in U)\bigr{\rangle}+\langle I(i\in U^{c})\rangle$
$=1-(1-\tanh(\beta M+h))\langle I(i\in U)\rangle.$
From the definition of $U$, it follows that w.h.p.,
$\sum_{(i,j)\in E}I(i\in U)$ $\geqslant\frac{1}{2}\sum_{i\in U}d_{i}\geqslant\frac{cdN}{8}.$
Therefore, we have w.h.p.,
$\frac{2}{Nd}\sum_{(i,j)\in E}\langle\sigma_{i}\sigma_{j}\rangle$ $\leqslant\frac{2}{Nd}\Bigl{(}|E|-(1-\tanh(\beta M+|h|))\Bigl{\langle}\sum_{(i,j)\in E}I(i\in U)\Bigr{\rangle}\Bigr{)}$
$\leqslant 1-(1-\tanh(\beta M))\frac{2}{Nd}\cdot\frac{cdN}{8}=1-(1-\tanh(\beta M))\frac{c}{4},$
which implies that there exists a constant $c^{\prime}>0$ such that
$\partial_{\beta}\operatorname{\mathds{E}}F_{z}(\beta,h)$ $\geqslant\beta^{2}c^{\prime}>0.$
Hence,
$\partial_{\beta}\operatorname{\mathds{E}}F_{g}(\beta,h)$ $\geqslant\beta^{2}c^{\prime}-\frac{C\beta^{3}}{\sqrt{d}}.$ (37)
Finally, from the concentration of the free energy in the Proposition 3.2 and the Lemma 3.1, we have
$\liminf_{N\to\infty}\bigl{(}F_{g}(\beta,h)-F_{g}(0,h)\bigr{)}$ $=\liminf_{N\to\infty}\bigl{(}\operatorname{\mathds{E}}F_{z}(\beta,h)-\operatorname{\mathds{E}}F_{z}(0,h)\bigr{)}$
$=\liminf_{N\to\infty}\int_{0}^{\beta}\partial_{t}\operatorname{\mathds{E}}F_{z}(t,h)dt$
$\geqslant\liminf_{N\to\infty}(\beta-\beta_{0})\partial_{t}\operatorname{\mathds{E}}F_{z}(\beta_{0},h)>0$
for $\beta_{0}<\beta,$ where the last inequality used the fact that $\partial_{t}\operatorname{\mathds{E}}F_{z}(t,h)$ is a nondecreasing function in $t$ as well as (37). The asserted exponential tail probability is guaranteed by Proposition 3.2.
$\blacksquare$
4. Proof of Theorem 1.5
We first establish the bound for the expected operator norm of $J=\frac{1}{\sqrt{d}}A\circ G$. It relies on the following fundamental result in [BvH16]*Corollary 3.3, which we recall as follows.
Lemma 4.1 ([BvH16]*Corollary 3.3).
@@
We will apply the result [CDG23]*Proposition A.1 for subgaussian random matrix to complete the proof. Note that in our setting, the operator norm is obviously a convex function. On the other hand, since for general matrices $A,B$,
@@
where $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm. This implies the Lipschitz property of the operator norm. There are in total $Nd$ many nonzero entries in $J$, then directly from [CDG23]*Proposition A.1,
@@
Using the Assumption 1.3, it is easy to see that $\operatorname{\mathds{E}}\left\|J\right\|=O(1)$ from Lemma 4.2. On the other hand, since $d\gg\log d$, the r.h.s in the concentration bound of Lemma 4.3 can be bounded as follows,
$2\exp\left(-\frac{Cdt^{2}}{K^{2}\log(Nd)}\right)\leqslant 2\exp\left(-\frac{Cdt^{2}}{K^{2}\log(N)}\right).$
@@
To establish this lemma, we need an elementary result that states that given a graph $\mathcal{G}$, we can always find a good expander subset $S\subseteq[N]$ with a large number of outgoing edges, see, e.g. [AS08]*Theorem 2.2.2.
@@
which contains the vertices in $S$ with at least ${d}/{4}$ many outgoing edges to $S^{c}$. Now let $I$ be a uniform random variable on $S$. Write
@@
As for the second probability, consider an arbitrary partition $\mathcal{D}=(e_{i})_{i\in[2M]}$ of $[-M,M]$ with $|e_{i}-e_{i-1}|<\varepsilon$ for $i\in[2M]$. From this,
@@
$\mathbb{P}\Bigl{(}\frac{1}{N}\min_{{\sigma}}\inf_{\gamma\in[-M,M]}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|\leqslant c_{1}\Bigr{)}\leqslant\mathbb{P}\Bigl{(}\frac{1}{N}\min_{{\sigma}}\inf_{\gamma\in\mathcal{D}}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|\leqslant c_{1}+\varepsilon^{2}\Bigr{)}.$
@@
$\mathbb{P}\Bigl{(}\min_{{\sigma}}\inf_{\gamma\in\mathcal{D}}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|\leqslant\delta\sqrt{N}\Bigr{)}\leqslant c^{\prime}\cdot 10^{-N}\cdot 2^{N}\cdot M,$
@@
Lemma 5.7 ([RV08]*Lemma 2.2).
Let $\zeta_{1},\ldots\zeta_{n}$ be independent random variables. Assume that there exist constants $K$ and $\varepsilon_{0}>0$ such that for each $i$, $\mathbb{P}(\left|\zeta_{i}\right|<\varepsilon)\leqslant K\varepsilon$ for all $\varepsilon\geqslant\varepsilon_{0}$. Then there exists an absolute constant $C_{1}$ depending only on $K$ and $\varepsilon_{0}$ such that
@@
Our goal is to combine Lemma 5.6 and Lemma 5.7 to establish some small-ball probability estimate for the vector ${m}({\sigma})$. However, because the matrix $J$ is symmetric, the random variables $m_{1}({\sigma}),m_{2}({\sigma}),\ldots,m_{N}({\sigma})$ are not independent. To circumvent this obstacle, we will make use of the fact that the entries in any off-diagonal block sub-matrices of $J$ are independent. In particular, we shall work with the sub-matrix $D_{2}$ of $J$ defined on the block $T\times T^{c}$, where $T$ is a $(1/(16C),d/4)$-good subset as in Lemma 5.2. With this choice of $T$, we partition the vertex set as $[N]=T\cup([N]\setminus T)$ and accordingly write ${\sigma}=({\sigma}^{(1)},{\sigma}^{(2)})\in\{-1,+1\}^{T}\times\{-1,+1\}^{[N]\setminus T}$. Similarly we represent $J$ (after possible permutation of the $N$ vertices) in the following block form,
Figure 1. Block decomposition of matrix $J$, where $D_{1}$ is a $\left|T\right|\times\left|T\right|$ matrix with row and column index set $T$, while $D_{2}$ has row index set $T$ and column index set $[N]\setminus T$. It can be seen that the entries in the block $D_{2}$ are independent from each other.
@@
$\sup_{v\in\mathbb{R}}\mathbb{P}\bigl{(}\bigl{|}(D_{2}{\sigma}^{(2)})_{i}-v\bigr{|}\leqslant\varepsilon\bigr{)}\lesssim\varepsilon.$
Combining with Lemma 5.7, there exists a constant $\delta>0$ such that for any $A\subseteq T$ with $\left|A\right|\geqslant\left|T\right|/{2}\geqslant{N}/(32C)$, we have
$\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\bigl{(}\bigl{\|}(D_{2}{\sigma}^{(2)})_{i\in A}-\boldsymbol{v}\bigr{\|}_{2}^{2}\leqslant\delta^{2}N\bigr{)}\lesssim 10^{-N}.$
@@
$\left\|{m}_{A}({\sigma})-\gamma\boldsymbol{1}\right\|_{2}^{2}\geqslant\bigl{\|}(D_{1}{\sigma}^{(1)}+D_{2}{\sigma}^{(2)})_{i\in A}-\gamma\boldsymbol{1}_{A}\bigr{\|}_{2}^{2}.$
@@
$\leqslant$ $\operatorname{\mathds{E}}_{D_{1}}\Bigl{[}\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\bigl{(}\bigl{\|}(D_{2}{\sigma}^{(2)}\bigr{\|}_{i\in A}-\boldsymbol{v}\bigr{\|}_{2}^{2}\leqslant\delta^{2}N\bigr{)}\Bigr{]}$
$\leqslant$ $\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\bigl{(}\bigl{\|}(D_{2}{\sigma}^{(2)})_{i\in A}-\boldsymbol{v}\bigr{\|}_{2}^{2}\leqslant\delta^{2}N\bigr{)}\leqslant\frac{K}{10^{N}}$
@@
$\left|B_{{\sigma}}^{+}\right|\sim\mathrm{Binomial}\left(\left|T\right|,p\right)\,\,\text{and}\,\,\left|B_{{\sigma}}^{-}\right|\sim\mathrm{Binomial}\left(\left|T\right|,p\right),$
where $p$ is some constant with $p\geqslant 1-{4C}/{M^{2}}$. Now, by the binomial concentration, one can choose sufficiently large $M$ such that
@@
where $\mathcal{D}$ is an arbitrary partition of $[-M,M]$ with mash $(\delta/2)$ and $|\mathcal{D}|\leqslant 10M\delta^{-1}$. Applying Proposition 5.4 for a fixed $\gamma$ and ${\sigma}$ in (41) and then taking a union bound over $\gamma\in\mathcal{D}$ and ${\sigma}$, we obtain, for $c_{0}=(c^{\prime\prime})^{2}\delta^{2}/2$,

### v2

@@ 2.1. Auxiliary Lemmas
We first establish the bound for the expected operator norm of $J=\frac{1}{\sqrt{d}}A\circ G$. It relies on the following fundamental result in Corollary 3.3 of [BvH16], which we recall as follows.
Lemma 3.1 (Corollary 3.3 [BvH16]).
@@
We will apply the Proposition A.1 in [CDG23] for subgaussian random matrix to complete the proof. Note that in our setting, the operator norm is obviously a convex function. On the other hand, since for general matrices $A,B$,
@@
where $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm. This implies the Lipschitz property of the operator norm. There are in total $Nd$ many nonzero entries in $J$, then directly from Proposition A.1 of [CDG23],
@@
Using the Assumption 1.3, it is easy to see that $\operatorname{\mathds{E}}\left\|J\right\|=O(1)$ from Lemma 3.2. On the other hand, since $d\gg\log N$, the r.h.s in the concentration bound of Lemma 3.3 can be bounded as follows,
$2\exp\left(-\frac{Cdt^{2}}{K^{2}\log(Nd)}\right)\leqslant 2\exp\left(-\frac{C^{\prime}dt^{2}}{K^{2}\log(N)}\right).$
@@
To establish this lemma, we need an elementary result that states that given a graph $\mathcal{G}$, we can always find a good expander subset $S\subseteq[N]$ with a large number of outgoing edges, see, e.g. Theorem 2.2.2 of [AS08].
@@
which contains the vertices in $S$ with at least d/4{d}/{4} many outgoing edges to $S^{c}$. Now let $I$ be a uniform random variable on $S$. Write
@@
As for the second probability, consider an arbitrary partition $\mathcal{D}=(e_{i})_{i\in[2M]}$ of [M,M][-M,M] with $|e_{i}-e_{i-1}|<\varepsilon$ for $i\in[2M]$. From this,
@@
$\mathbb{P}\Bigl{(}\frac{1}{N}\min_{{\sigma}}\inf_{\gamma\in[-M,M]}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|_{2}\leqslant c_{1}\Bigr{)}\leqslant\mathbb{P}\Bigl{(}\frac{1}{N}\min_{{\sigma}}\inf_{\gamma\in\mathcal{D}}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|_{2}\leqslant c_{1}+\varepsilon^{2}\Bigr{)}.$
@@
$\mathbb{P}\Bigl{(}\min_{{\sigma}}\inf_{\gamma\in\mathcal{D}}\left\|{m}({\sigma})-\gamma\boldsymbol{1}\right\|_{2}\leqslant\delta\sqrt{N}\Bigr{)}\leqslant c^{\prime}\cdot 10^{-N}\cdot 2^{N}\cdot M,$
@@
Lemma 4.7 (Lemma 2.2 [RV08]).
Let $\zeta_{1},\ldots\zeta_{n}$ be independent random variables. Assume that there exist constants $K$ and $\varepsilon_{0}>0$ such that for each $i$, $\mathbb{P}(\left|\zeta_{i}\right|<\varepsilon)\leqslant K\varepsilon$ for all $\varepsilon\geqslant\varepsilon_{0}$. Then there exists an absolute constant $C_{1}$ such that
@@
Our goal is to combine Lemma 4.6 and Lemma 4.7 to establish some small-ball probability estimate for the vector ${m}({\sigma})$. However, because the matrix $J$ is symmetric, the random variables $m_{1}({\sigma}),m_{2}({\sigma}),\ldots,m_{N}({\sigma})$ are not independent. To circumvent this obstacle, we will make use of the fact that the entries in any off-diagonal block sub-matrices of $J$ are independent. In particular, we shall work with the sub-matrix $D_{2}$ of $J$ defined on the block $T\times T^{c}$, where $T$ is a $(1/(16C),d/4)$-good subset as in Lemma 4.2. With this choice of $T$, we partition the vertex set as $N]=T\cup([N]\setminus T)$ and accordingly write ${\sigma}=({\sigma}^{(1)},{\sigma}^{(2)})\in\{-1,+1\}^{T}\times\{-1,+1\}^{[N]\setminus T}$. Similarly we represent $J$ (after possible permutation of the $N$ vertices) in the following block form,
Figure 2. Block decomposition of matrix $J$, where $D_{1}$ is a $\left|T\right|\times\left|T\right|$ matrix with row and column index set $T$, while $D_{2}$ has row index set $T$ and column index set $N]\setminus T$. It can be seen that the entries in the block $D_{2}$ are independent from each other.
@@
$\sup_{v\in\mathbb{R}}\mathbb{P}\bigl{(}\bigl{|}(D_{2}{\sigma}^{(2)})_{i}-v\bigr{|}\leqslant\varepsilon\bigr{)}\leqslant 2\sqrt{\frac{2}{\pi}}\varepsilon+C_{0}B\left(\frac{2}{\sqrt{d}}\right)^{1/3}\leqslant K^{\prime}(\varepsilon+d^{-1/6}),$ (36)
where $K^{\prime}$ is a constant only depending on $B$. Note that the hypothesis of Lemma 4.7 holds for $\varepsilon_{0}=d^{-1/6}$ with $K=2K^{\prime}$. Therefore, for any constant $\delta>0$ and for $d$ sufficiently large such that $\varepsilon_{0}=d^{-1/6}\leqslant\delta$, by Lemma 4.7, we obtain for any $A\subseteq T$ with $\left|A\right|\geqslant\left|T\right|/{2}\geqslant{N}/(32C)$,
$\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\bigl{(}\bigl{\|}(D_{2}{\sigma}^{(2)})_{i\in A}-\boldsymbol{v}\bigr{\|}_{2}^{2}\leqslant\delta^{2}N\bigr{)}\leqslant(C_{1}K\delta)^{N/(32C)}\lesssim 10^{-N},$
where $C_{1}$ is as given in the Lemma 4.7 and the last inequality follows by taking $\delta>0$ sufficiently small constant.
@@
$\left\|{m}_{A}({\sigma})-\gamma\boldsymbol{1}_{A}\right\|_{2}^{2}\geqslant\bigl{\|}(D_{1}{\sigma}^{(1)}+D_{2}{\sigma}^{(2)})_{i\in A}-\gamma\boldsymbol{1}_{A}\bigr{\|}_{2}^{2}.$
@@
$\leqslant$ $\operatorname{\mathds{E}}_{D_{1}}\Bigl{[}\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\left(\left\|(D_{2}{\sigma}^{(2)})_{i\in A}-\boldsymbol{v}\right\|_{2}^{2}\leqslant\delta^{2}N\right)\Bigr{]}$
$\leqslant$ $\sup_{\boldsymbol{v}\in\mathbb{R}^{\left|A\right|}}\mathbb{P}\left(\left\|(D_{2}{\sigma}^{(2)})_{i\in A}-\boldsymbol{v}\right\|_{2}^{2}\leqslant\delta^{2}N\right)\leqslant\frac{K}{10^{N}}$
@@
$\left|B_{{\sigma}}^{+}\right|\sim\mathrm{Binomial}\left(\left|T\right|,p_{+}\right)\,\,\text{and}\,\,\left|B_{{\sigma}}^{-}\right|\sim\mathrm{Binomial}\left(\left|T\right|,p_{-}\right),$
where $p_{-},p_{+}$ are some constants greater than $1-{4C}/{M^{2}}$. Now, by the binomial concentration, one can choose sufficiently large $M$ such that
@@
where $\mathcal{D}$ is an arbitrary partition of [M,M][-M,M] with mash $(\delta/2)$ and $|\mathcal{D}|\leqslant 10M\delta^{-1}$. Applying Proposition 4.4 for a fixed $\gamma$ and $\sigma$} in (37) and then taking a union bound over $\gamma\in\mathcal{D}$ and $\sigma$}, we obtain, for $c_{0}=(c^{\prime\prime})^{2}\delta^{2}/2$,